### v2

@@ 1 Introduction
Modeling user choices in this setting faces several challenges that are present on many other online platforms. First, conditional on subscribing to the platform, users costlessly consume goods, meaning that there is no price variation that we can exploit to help us measure user valuation of goods or ascertain the role of recommendations. Second, estimating movie demand has historically been challenging as traditional good characteristics, such as cast and genre, have done a relatively poor job of predicting user demand (Einav, 2007; McKenzie, 2023). Finally, there is potential serial dependence in choices (i.e., the good a user consumes today influences their choice tomorrow), and both the number of goods and users is incredibly large. We overcome each of these challenges and provide a model that we can apply to estimate the value of goods and recommendations.
The structure of our model follows a low-rank discrete choice model (Kallus and Udell, 2016), which assumes that the good $\times$ user matrix of user preferences can be represented by a low-rank representation of good and user embeddings. We extend this baseline model in several ways to solve the aforementioned challenges. For the first challenge, we model the effects of recommendation via an additive recommendation “bonus” that provides a utility boost for goods that are recommended. This is a reduced-form characterization of the effects of the recommendation as reducing information frictions for these goods (Aridor et al., 2023). For the second challenge, we endogenously learn the good-level “characteristics” directly from user choices, circumventing the need to define a reasonable set of characteristics that can rationalize choices. For the third challenge, we adapt a popular approach in the machine learning literature to our setting and consider a transformer-style architecture (Vaswani et al., 2017) that extends standard state-dependent demand methods to allow for user preferences to flexibly adapt across time and rely on a user’s consumption history. [^1] The central concept of the transformer architecture is “attention”, which is a context-dependent weight computed for each vector in a sequence, enabling a weighted-average summarization of an arbitrary-length sequence into a single vector. In this work, we use a simpler form of attention weighting described in Section 3.
@@
A key aspect of our approach and its validation is to rely on idiosyncratic randomization that the recommender system (RecSys) induces to learn user preferences, which allows us to separate out the effects of recommendation from the “intrinsic” preferences of users. This randomization occasionally induces variation in the selection and ordering of goods in the recommendations. To validate our model, we run an experiment that exogenously “nudges” the salience of a random set of goods, allowing us to generate model-free diversion ratio estimates for this set of goods (Conlon and Mortimer, 2021). We generate the same diversion ratio estimates using our model and find that, out of sample, the model’s predicted diversion has a 0.86 correlation and 0.73 $R^{2}$ when compared to the model-free diversion ratios. This approach builds upon a recent line of work in industrial organization that relies on exogenous non-price variation, such as good availability experiments, in order to estimate diversion ratios in settings without price variation (Conlon and Mortimer, 2013; Conlon et al., 2023; Aridor, 2025).
@@
In order to quantify the overall value of the RecSys, we evaluate counterfactuals where the set of recommendations is generated by (i) randomized, (ii) popularity-based, and (iii) traditional matrix factorization (Koren et al., 2009) recommendation systems. [^2] As discussed in Gomez-Uribe and Hunt (2015), a matrix factorization based RecSys was the algorithm used at Netflix 10 years prior to this work. Our main finding is that the current RecSys balances high engagement with high consumption diversity. Moving to an entirely random recommendation policy decreases engagement by 16%, with minimal improvements to overall consumption diversity. Moving to a popularity-based policy or a traditional matrix factorization approach decreases engagement by 12% and 4%, respectively, while having large negative effects on consumption diversity. Internal estimates of the value of engagement suggest that these are economically large changes (cf. Gomez-Uribe and Hunt, 2015). Furthermore, the current RecSys dramatically improves consumption diversity relative to the latter benchmarks, which is important for long-run user satisfaction (Anderson et al., 2020) and a healthy two-sided market that distributes viewership more evenly across good producers.
@@
The Impact of Online Recommendations. We contribute to a burgeoning literature that quantifies the effects of recommendations and rankings on choices. Senecal and Nantel (2004), Das et al. (2007), Freyne et al. (2009), Zhou et al. (2010), Peukert et al. (2024), Holtz et al. (2020), Aridor et al. (2023), and Donnelly et al. (2024) show that, compared to non-personalized benchmarks, recommendation systems increased consumption in hypothetical choices in a lab experiment, Google News, a social network, a news website, YouTube, Spotify, MovieLens, and Wayfair, respectively. These papers collectively provide evidence that personalized recommendations, relative to a non-personalized benchmark, meaningfully increase consumption. There has subsequently been a significant amount of work measuring how personalization impacts both the user-level and aggregate-level diversity of the consumption distribution (Fleder and Hosanagar, 2009; Nguyen et al., 2014; Brynjolfsson et al., 2011; Hosanagar et al., 2013; Lee and Hosanagar, 2019; Holtz et al., 2020; Korganbekova and Zuber, 2023; Calvano et al., 2022).
Our work contributes to the literature in several ways. First, we quantify the incremental value of recommendation on Netflix, a platform with a highly influential and economically significant RecSys (Bennett and Lanning, 2007). By comparing not only to non-personalized benchmarks, but also a previous state of the art algorithm (e.g., matrix factorization (Koren et al., 2009)), we quantify the value of advances in personalization technology in the past decade (e.g., relative to those discussed in Gomez-Uribe and Hunt (2015)). We showcase how this leads to differences in overall engagement as well as its influence on consumption diversity, which is especially important in two-sided markets. Second, by systematically quantifying the value of recommendation across the good distribution and isolating how much of this comes from effective targeting, we more comprehensively document the heterogeneity in the value of recommendations across different goods—highlighting that mid-sized goods benefit the most and not goods with broad appeal or very niche goods (in contrast to the hypothesis of Anderson et al. (2006)). Third, similar to the insights of Morozov et al. (2021) in other markets with search and information frictions, we highlight the importance of accounting for the influence of recommendations when estimating user preferences.
@@
First, we contribute to the growing literature on measuring substitution patterns in markets with limited price variation. While these methods are broadly applicable, they are particularly relevant in the “attention economy,” where the scarce resource is user attention and goods are typically costless to consume (Brynjolfsson et al., 2019; Calvano and Polo, 2021; Yuan, 2025). Existing work relies on generated or exogenous good unavailability variation (Conlon and Mortimer, 2013; Raval et al., 2022; Aridor, 2025) or using survey-based methods to elicit choices under hypothetical scenarios (Dertwinkel-Kalt et al., 2024; Bursztyn et al., 2025) to estimate second-choice diversion ratios. Facing a similar challenge, we highlight that platforms can use exploration experiments in the RecSys as plausibly exogenous variation to estimate both choice models and model-free diversion ratios (Conlon and Mortimer, 2021).
Second, we contribute to the literature on state-dependent demand estimation, which provides a class of models to rationalize persistence in user choices. The empirical challenge in this literature, dating back to Heckman (1981, 1991), is to disentangle whether past choices causally affect future choices or if choices are spuriously serially correlated due to unobserved preference heterogeneity. The canonical models of state-dependent demand (Dubé et al., 2010; Simonov et al., 2020) provide a Markovian formulation with rich preference heterogeneity to solve this problem. We formulate our problem in a similar manner, except in our context we are more interested in how to model the evolution of preferences over time as opposed to decoupling these two channels. In order to do so, adapt methods from transformer-style architectures that are popular in machine learning (Vaswani et al., 2017) to characterize state-dependence based on the full consumption history. This allows users’ tastes to flexibly change over time and computationally simplifies our model by requiring us to only keep track of different consumption histories, instead of the full matrix of user embeddings. The latter enables us to estimate this model at the scale of millions of users with minimal compute requirements. Topically, Zeller (2022); Lee et al. (2025) apply models with state dependence in media streaming contexts, albeit with substantively different models than we consider here.
@@ 2 Institutional Details
To help users find satisfying goods, Netflix’s homepage is personalized. There has been considerable development on the recommendation system, dating back to the public Netflix prize (Bennett and Lanning, 2007) and disclosed in Gomez-Uribe and Hunt (2015). The implementation discussed in Gomez-Uribe and Hunt (2015) relies on matrix factorization approaches, but since then the RecSys has evolved to incorporate newer algorithmic advances in generating recommendations (e.g., see Steck et al. (2021), Joshi et al. (2024)).
@@ 3 Demand Model
This utility formulation is intrinsically in good space as it considers that our goal is to learn the matrix of user $\times$ good utilities. We follow a similar formulation as Kallus and Udell (2016) by assuming that this can be decomposed into low-rank good and user matrices. We exogenously impose the rank of these matrices, such that $B$ is $\mathbb{R}^{J\times d}$ and $A_{t}$ is $\mathbb{R}^{I\times d}$. Beyond these two terms, which represent the “intrinsic” preferences that users have for the goods, we explicitly model the role of recommendations where $\mathcal{C}_{it}\equiv\cup_{r\in\{\text{billboard, top 25, top 100}\}}\mathcal{C}_{irt}$ denotes the set of recommendations that user $i$ observes at time $t$. We handle the granularity of the different types of recommendations (as described in Section 2) by having separate recommendation utility booster terms for the billboard, top 25, and top 100 recommendations (denoted by $\beta_{jr}$). Finally, $\varepsilon_{ijt}$ denotes the standard Type-1 Extreme Value error. Under the standard assumption that $\varepsilon_{ijt}$ are independent and identically distributed, this induces the probability that good $j$ will be chosen by user $i$ in time period $t$:
@@
The first challenge is that, in our empirical context, recommendations play a large role in driving user choices. In principle, recommendations can influence user choices through two mechanisms: making users aware of goods they did not know existed and alleviating information or search frictions conditional on awareness. We follow Aridor et al. (2023) in assuming that the latter is the dominant force and, as such, model the effect of recommendation as providing an additive utility bonus for the set of goods that are recommended, which provides a reduced-form formulation of the (positive) informational role that recommendations play in helping guide user choices.
The second challenge is that modeling demand for movies and TV shows has traditionally been difficult since traditional good characteristics (e.g., genre, cast, etc.) are typically not sufficient to explain user choices (Basuroy et al., 2003; Einav, 2007; McKenzie, 2023). Instead of relying on traditional characteristics, we endogenously learn the good embeddings via user choices in a similar spirit as traditional matrix factorization approaches used in recommender systems (Mnih and Salakhutdinov, 2007; Koren et al., 2009). As in Kallus and Udell (2016), the key difference is that we embed this within a discrete-choice model that allows us to flexibly learn these good embeddings. While this is our preferred specification, we nonetheless show that the model performs reasonably well if we use good-level embeddings that are learned on pre-consumption good characteristics and tags in Appendix B.
@@
The third challenge is that there may be serial dependence in choices, for instance due to users being more likely to watch more sports-related show after watching a sports-related movie. We let the preference weights, $A_{it}$, vary over time and encapsulate everything the user $i$ has watched up to period t1t-1, based on a sequence model of prior watch history as diagrammed in Figure 2. At a given time $t$, we encode the good embeddings $B_{j}$ in an ordered sequence of when the user watched them. We pass that vector through a two-layer fully-connected MLP that produces, for each time period $t$, a set of attention weights for each previous time period and a transformed, user-specific, set of characteristics. Taking the weighted sum of these and passing it through a two-layer fully connected MLP again provides our vector of estimates $A_{it}$. Intuitively, the formulation is similar to a traditional sequence model in machine learning where the future goods that a user consumes depend on their previous sequence and the model endogenously learns position-specific attention weights. This aspect of the formulation is a modified variant of state-dependent demand models (Dubé et al., 2010) that combines both time-variation in preferences and “intrinsic” preferences. In addition to addressing serial dependence, this architecture avoids the need to learn user embeddings directly, which becomes highly computationally challenging at the scale of millions of users.
Identification. There are several identification challenges that we face. The first is that recommendations are endogenously targeted towards users because the recommender system predicts that they have high predicted utility for that user. We rely on the intrinsic randomization introduced into the recommender system for the purposes of exploration as exogenous variation in the set of recommendations that breaks this endogenous targeting. This exogenous variation also allows us to identify the effect of state-dependence since users are more likely to consume goods that are present in the recommendations and thus the exogenous variation in choice probabilities provides us with randomization in the initial set of consumed goods. Additionally, the same identification argument as Zeller (2022) applies in our context since goods have a finite amount of consumption time.
@@ 4.1 Model Estimates
In Figure 4 we show how a projection of the good embeddings via UMAP (McInnes et al., 2018) corresponds to characteristic good attributes. Figure 4(a) shows that goods that are distinctly different from others (e.g., kids and documentaries content) are clustered together, while the remaining genres are similarly clustered together though there is considerable overlap with other genres. [^4] This highlights the challenges with using observable characteristics, such as cast and genre, alone. Additionally, Figure 4(b) shows a distinct separation between non-English and English goods and Figure 4(c) highlights that, within clusters, there is a clear separation between movies and television shows. Overall, these figures highlight that the endogenously learned characteristic space is reasonable and accords with intuition, but also subtly captures more complex similarity between goods than can be discerned from standard observable characteristics.

### v3

@@ 1 Introduction
Modeling user choices in this setting faces several challenges that are present on many other online platforms. First, conditional on subscribing to the platform, users costlessly consume goods, meaning that there is no price variation that we can exploit to help us measure user valuation of goods or ascertain the role of recommendations. Second, estimating movie demand has historically been challenging as traditional good characteristics, such as cast and genre, have done a relatively poor job of predicting user demand (einav2007seasonality; mckenzie2023economics). Finally, there is potential serial dependence in choices (i.e., the good a user consumes today influences their choice tomorrow), and both the number of goods and users is incredibly large. We overcome each of these challenges and provide a model that we can apply to estimate the value of goods and recommendations.
The structure of our model follows a low-rank discrete choice model (kallus2016revealed), which assumes that the good $\times$ user matrix of user preferences can be represented by a low-rank representation of good and user embeddings. We extend this baseline model in several ways to solve the aforementioned challenges. For the first challenge, we model the effects of recommendation via an additive recommendation “bonus” that provides a utility boost for goods that are recommended. This is a reduced-form characterization of the effects of the recommendation as reducing information frictions for these goods (aridor2022economics). For the second challenge, we endogenously learn the good-level “characteristics” directly from user choices, circumventing the need to define a reasonable set of characteristics that can rationalize choices. For the third challenge, we adapt a popular approach in the machine learning literature to our setting and consider a transformer-style architecture (vaswani2017attention) that extends standard state-dependent demand methods to allow for user preferences to flexibly adapt across time and rely on a user’s consumption history. [^1] The central concept of the transformer architecture is “attention”, which is a context-dependent weight computed for each vector in a sequence, enabling a weighted-average summarization of an arbitrary-length sequence into a single vector. In this work, we use a simpler form of attention weighting described in Section 3.
@@
A key aspect of our approach and its validation is to rely on idiosyncratic randomization that the recommender system (RecSys) induces to learn user preferences, which allows us to separate out the effects of recommendation from the “intrinsic” preferences of users. This randomization occasionally induces variation in the selection and ordering of goods in the recommendations. To validate our model, we run an experiment that exogenously “nudges” the salience of a random set of goods, allowing us to generate model-free diversion ratio estimates for this set of goods (conlon2021empirical). We generate the same diversion ratio estimates using our model and find that, out of sample, the model’s predicted diversion has a 0.86 correlation and 0.73 $R^{2}$ when compared to the model-free diversion ratios. This approach builds upon a recent line of work in industrial organization that relies on exogenous non-price variation, such as good availability experiments, in order to estimate diversion ratios in settings without price variation (conlon2013experimental; conlon2023estimating; aridor2025measuring).
@@
In order to quantify the overall value of the RecSys, we evaluate counterfactuals where the set of recommendations is generated by (i) randomized, (ii) popularity-based, and (iii) traditional matrix factorization (koren2009matrix) recommendation systems. [^2] As discussed in gomez2015netflix, a matrix factorization based RecSys was the algorithm used at Netflix 10 years prior to this work. Our main finding is that the current RecSys balances high engagement with high consumption diversity. Moving to an entirely random recommendation policy decreases engagement by 16%, with minimal improvements to overall consumption diversity. Moving to a popularity-based policy or a traditional matrix factorization approach decreases engagement by 12% and 4%, respectively, while having large negative effects on consumption diversity. Internal estimates of the value of engagement suggest that these are economically large changes (cf. gomez2015netflix). Furthermore, the current RecSys dramatically improves consumption diversity relative to the latter benchmarks, which is important for long-run user satisfaction (anderson2020algorithmic).
@@
The Impact of Online Recommendations. We contribute to a burgeoning literature that quantifies the effects of recommendations and rankings on choices. SenecalNantel2004, DasDatarGargRajaram2007, FreyneJacoviGuyGeyer2009, ZhouKhemmaratGao2010, claussen2019editor, HoltzCarteretteChandarNazariCramerAral2020WP, aridor2022economics, and donnelly2024welfare show that, compared to non-personalized benchmarks, recommendation systems increased consumption in hypothetical choices in a lab experiment, Google News, a social network, a news website, YouTube, Spotify, MovieLens, and Wayfair, respectively. These papers collectively provide evidence that personalized recommendations, relative to a non-personalized benchmark, meaningfully increase consumption. There has subsequently been a significant amount of work measuring how personalization impacts both the user-level and aggregate-level diversity of the consumption distribution (FlederHosanagar2009MS; NguyenHuiHarperTerveenKonstan2014; BrynjolfssonHuSimester2011MS; HosanagarFlederLeeBuja2013MS; LeeHosanagar2019ISR; HoltzCarteretteChandarNazariCramerAral2020WP; KorganbekovaZuber2023WP; CalvanoCalzolariDenicoloPastorello2022WP).
Our work contributes to the literature in several ways. First, we quantify the incremental value of recommendation on Netflix, a platform with a highly influential and economically significant RecSys (bennett2007netflix). By comparing not only to non-personalized benchmarks, but also a previous state of the art algorithm (e.g., matrix factorization (koren2009matrix)), we quantify the value of advances in personalization technology in the past decade (e.g., relative to those discussed in gomez2015netflix). We showcase how this leads to differences in overall engagement as well as its influence on consumption diversity, which is especially important in large catalogs. Second, by systematically quantifying the value of recommendation across the good distribution and isolating how much of this comes from effective targeting, we more comprehensively document the heterogeneity in the value of recommendations across different goods—highlighting that mid-sized goods benefit the most and not goods with broad appeal or very niche goods (in contrast to the hypothesis of anderson2006long). Third, similar to the insights of morozov2021estimation in other markets with search and information frictions, we highlight the importance of accounting for the influence of recommendations when estimating user preferences.
@@
First, we contribute to the growing literature on measuring substitution patterns in markets with limited price variation. While these methods are broadly applicable, they are particularly relevant in the “attention economy,” where the scarce resource is user attention and goods are typically costless to consume (brynjolfsson2019using; calvano2021market; yuan2025competing). Existing work relies on generated or exogenous good unavailability variation (conlon2013experimental; raval2022using; aridor2025measuring) or using survey-based methods to elicit choices under hypothetical scenarios (dertwinkel2024defining; bursztyn2025measuring) to estimate second-choice diversion ratios. Facing a similar challenge, we highlight that platforms can use exploration experiments in the RecSys as plausibly exogenous variation to estimate both choice models and model-free diversion ratios (conlon2021empirical).
Second, we contribute to the literature on state-dependent demand estimation, which provides a class of models to rationalize persistence in user choices. The empirical challenge in this literature, dating back to heckman1981heterogeneity; heckman1991identifying, is to disentangle whether past choices causally affect future choices or if choices are spuriously serially correlated due to unobserved preference heterogeneity. The canonical models of state-dependent demand (dube2010state; simonov2020state) provide a Markovian formulation with rich preference heterogeneity to solve this problem. We formulate our problem in a similar manner, except in our context we are more interested in how to model the evolution of preferences over time as opposed to decoupling these two channels. In order to do so, adapt methods from transformer-style architectures that are popular in machine learning (vaswani2017attention) to characterize state-dependence based on the full consumption history. This allows users’ tastes to flexibly change over time and computationally simplifies our model by requiring us to only keep track of different consumption histories, instead of the full matrix of user embeddings. The latter enables us to estimate this model at the scale of millions of users with minimal compute requirements. Topically, zeller20222drip; lee2025pay apply models with state dependence in media streaming contexts, albeit with substantively different models than we consider here.
@@ 2 Institutional Details
To help users find satisfying goods, Netflix’s homepage is personalized. There has been considerable development on the recommendation system, dating back to the public Netflix prize (bennett2007netflix) and disclosed in gomez2015netflix. The implementation discussed in gomez2015netflix relies on matrix factorization approaches, but since then the RecSys has evolved to incorporate newer algorithmic advances in generating recommendations (e.g., see steck2021deep, Lamkhede2024, hsiao2025foundation).
@@ 3 Demand Model
This utility formulation is intrinsically in good space as it considers that our goal is to learn the matrix of user $\times$ good utilities. We follow a similar formulation as kallus2016revealed by assuming that this can be decomposed into low-rank good and user matrices. We exogenously impose the rank of these matrices, such that $B$ is $\mathbb{R}^{J\times d}$ and $A_{t}$ is $\mathbb{R}^{I\times d}$. Beyond these two terms, which represent the “intrinsic” preferences that users have for the goods, we explicitly model the role of recommendations where $\mathcal{C}_{it}\equiv\cup_{r\in\{\text{billboard, top 25, top 100}\}}\mathcal{C}_{irt}$ denotes the set of recommendations that user $i$ observes at time $t$. We handle the granularity of the different types of recommendations (as described in Section 2) by having separate recommendation utility booster terms for the billboard, top 25, and top 100 recommendations (denoted by $\beta_{jr}$). Finally, $\varepsilon_{ijt}$ denotes the standard Type-1 Extreme Value error. Under the standard assumption that $\varepsilon_{ijt}$ are independent and identically distributed, this induces the probability that good $j$ will be chosen by user $i$ in time period $t$:
@@
The first challenge is that, in our empirical context, recommendations play a large role in driving user choices. In principle, recommendations can influence user choices through two mechanisms: making users aware of goods they did not know existed and alleviating information or search frictions conditional on awareness. We follow aridor2022economics in assuming that the latter is the dominant force and, as such, model the effect of recommendation as providing an additive (ex ante) utility bonus for the set of goods that are recommended, which provides a reduced-form formulation of the (positive) informational role that recommendations play in helping guide user choices.
The second challenge is that modeling demand for movies and TV shows has traditionally been difficult since traditional good characteristics (e.g., genre, cast, etc.) are typically not sufficient to explain user choices (basuroy2003critical; einav2007seasonality; mckenzie2023economics). Instead of relying on traditional characteristics, we endogenously learn the good embeddings via user choices in a similar spirit as traditional matrix factorization approaches used in recommender systems (mnih2007probabilistic; koren2009matrix). As in kallus2016revealed, the key difference is that we embed this within a discrete-choice model that allows us to flexibly learn these good embeddings. While this is our preferred specification, we nonetheless show that the model performs reasonably well if we use good-level embeddings that are learned on pre-consumption good characteristics and tags in Appendix B.
@@
The third challenge is that there may be serial dependence in choices, for instance due to users being more likely to watch more sports-related show after watching a sports-related movie. We let the preference weights, $A_{it}$, vary over time and encapsulate everything the user $i$ has watched up to period t1t-1, based on a sequence model of prior watch history as diagrammed in Figure 2. At a given time $t$, we encode the good embeddings $B_{j}$ in an ordered sequence of when the user watched them. We pass that vector through a two-layer fully-connected MLP that produces, for each time period $t$, a set of attention weights for each previous time period and a transformed, user-specific, set of characteristics. Taking the weighted sum of these and passing it through a two-layer fully connected MLP again provides our vector of estimates $A_{it}$. Intuitively, the formulation is similar to a traditional sequence model in machine learning where the future goods that a user consumes depend on their previous sequence and the model endogenously learns position-specific attention weights. This aspect of the formulation is a modified variant of state-dependent demand models (dube2010state) that combines both time-variation in preferences and “intrinsic” preferences. In addition to addressing serial dependence, this architecture avoids the need to learn user embeddings directly, which becomes highly computationally challenging at the scale of millions of users.
Identification. There are several identification challenges that we face. The first is that recommendations are endogenously targeted towards users because the recommender system predicts that they have high predicted utility for that user. We rely on the intrinsic randomization introduced into the recommender system for the purposes of exploration as exogenous variation in the set of recommendations that breaks this endogenous targeting. This exogenous variation also allows us to identify the effect of state-dependence since users are more likely to consume goods that are present in the recommendations and thus the exogenous variation in choice probabilities provides us with randomization in the initial set of consumed goods. Additionally, the same identification argument as zeller20222drip applies in our context since goods have a finite amount of consumption time.
@@ 4.1 Model Estimates
In Figure 4 we show how a projection of the good embeddings via UMAP (mcinnes2018umap-software) corresponds to characteristic good attributes. Figure 4(a) shows that goods that are distinctly different from others (e.g., kids and documentaries content) are clustered together, while the remaining genres are similarly clustered together though there is considerable overlap with other genres. [^4] This highlights the challenges with using observable characteristics, such as cast and genre, alone. Additionally, Figure 4(b) shows a distinct separation between non-English and English goods and Figure 4(c) highlights that, within clusters, there is a clear separation between movies and television shows. Overall, these figures highlight that the endogenously learned characteristic space is reasonable and accords with intuition, but also subtly captures more complex similarity between goods than can be discerned from standard observable characteristics.